how do you solve for x in this equation .6/256x^5=1.1*10^13 I know that x should equal 7.3*10^-4, but i do not know how to get that. Does anyone know the steps to solve this?
Answer by stanbon(75887)

solve for x in this equation .6/256x^5=1.1*10^13 I know that x should equal 7.3*10^-4,
-------------
0.6/256x^5 = 1.1*10^13
(2.34375*10^-3)x^5 = 1.1*10^13
---
x^5 = 1.1*10^13 / (2.34375*10^-3)
---
x^5 = 4.69333*10^15
---
x = 1.36237*10^3
==========================
I can't imagine where you get 7.3*10^-4 as the answer.
==========================
Cheers,
Stan H.